OpenStudy (anonymous):

The length of a rectangular table is given by 3x+5 cm and width is given by 2x-7 cm. Express the area of the rectangle in terms of x.

OpenStudy (anonymous):

(a+b)*(c+d)= a*c+b*c+a*d+b*d where a=3x b=5 c=2x d= -7

OpenStudy (lgbasallote):

area = length x width therefore \[A = (3x+5)(2x-7)\] expand this and you'll get your answer do you need more help?

OpenStudy (anonymous):

ahmmm.. thank you.. really help it a lot..

OpenStudy (lgbasallote):

welcome :D

OpenStudy (anonymous):

The base of a triangle measures 3y+7 inches. Express the area of the triangle in terms of y..

OpenStudy (anonymous):

area of a triangle is usually 1/2 base*height could you draw a picture

OpenStudy (lgbasallote):

not enough info

OpenStudy (anonymous):

The base of a triangle measures 3y+7 inches. Express the area of the triangle in terms of y..

OpenStudy (anonymous):

is it an equilateral triangle?

OpenStudy (anonymous):

The length and width of a rectangle are given by 2 consecutive odd integers. Write the expression of the area of the rectangle.

OpenStudy (anonymous):

area of a rectangle is length*width let x be an odd number and the length then the next consecutive odd integer would be x+2 and this would be the width so just multiply the length and width

OpenStudy (anonymous):

i cant understand..

OpenStudy (anonymous):

lets say x = 3 the next consecutive odd number would be 5 so what must we do to x in order to get 5?

OpenStudy (anonymous):

what is the first number?

OpenStudy (anonymous):

x=3 x+b=5 what is b

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

or 3 because its odd?

OpenStudy (anonymous):

so instead of x=3 we have x= an odd number right?

OpenStudy (anonymous):

the next consecutive odd number would be the original odd number +2 so (x+2)= next consecutive odd number if x is an odd number

OpenStudy (anonymous):

so what number should i apply in the problem?

OpenStudy (anonymous):

you shouldnt apply a number for the problem if it is not given, instead you use a variable like x, or y, or m, etc

OpenStudy (anonymous):

so x= length (x+2)= width Area of a rectangle= length*width Area= x*(x+2) distribute x state that x must be an odd number

OpenStudy (anonymous):

but in my problem the length and width in unknown?

OpenStudy (anonymous):

you're told that the length and width are 2 consecutive odd integers

OpenStudy (anonymous):

how can i get that?

OpenStudy (anonymous):

please give me an illustration..

OpenStudy (anonymous):